The area of a triangular poster is 84 square feet. Its height is 2 feet
shorter than the length of its base.
The formula for the area of a triangle is A = 1/2bh.
Find the height of the triangle.

Answer by nerdybill(7384)

The area of a triangular poster is 84 square feet. Its height is 2 feet
shorter than the length of its base.
The formula for the area of a triangle is A = 1/2bh.
Find the height of the triangle.
.
Let b = base
then
b-2 = height
.
we derive our equation from: A = (1/2)bh
84 = (1/2)b(b-2)
multiplying both sides by 2:
168 = b(b-2)
168 = b^2-2b
0 = b^2-2b-168
0 = (b-14)(b+12)
b = {-12, 14}
throw out the negative solution leaving
b = 14 feet (base)
.
height:
b-2 = 14-2 = 12 feet